OpenStudy (anonymous):

4) What is the height of the apple when your friend catches it? (**HINT: The height would be the y-coordinate of the highest point.) Double click on the space below to show your work. h(t) = -16t^2 + 38.4t + 0.96 h= height t= time (seconds) I will fan and medal for work plus answers!

OpenStudy (anonymous):

How do I find the height? How do I switch the equation to find the height instead of the time? I can solve it just not sure how to set it up.

OpenStudy (anonymous):

there is more to this question right?

OpenStudy (anonymous):

like more words

OpenStudy (anonymous):

Make t=0. :)

OpenStudy (anonymous):

The information is enough. :)

OpenStudy (anonymous):

naw dude its asking for the vertex

OpenStudy (anonymous):

Would I just substitute the time that we just figured out into the equation? The 2.4 seconds?

OpenStudy (anonymous):



OpenStudy (anonymous):

No you don't have to. Just make t=0. :) \(h(t) = -16(0)2 + 38.4(0) + 0.96\)

OpenStudy (anonymous):

why would you do that?? is the friend .96 ft off the ground

OpenStudy (anonymous):

The height we're finding is the same thing as the vertex. =_= Notice that the vertex in this case is the maximum because \(a\) is a negative.

OpenStudy (anonymous):

no not for t=0

OpenStudy (anonymous):

t=o is your initial starting point not the vertex

OpenStudy (anonymous):

or your y-intercept

OpenStudy (anonymous):

If (0,y) <--- this is the vertex. Please stop misguiding the one who asked.

OpenStudy (anonymous):

NO IT ISNT...

OpenStudy (anonymous):

If you solve the equation mathway gave me, I got 0.96.

OpenStudy (anonymous):

b/2a is the x coordinate and f(b/2a) is your Y

OpenStudy (anonymous):

"(**HINT: The height would be the y-coordinate of the highest point.)" Which means that you need to find the y-coordinate.

OpenStudy (anonymous):

Yes, it is \(\huge 0.96\). :)

OpenStudy (anonymous):

MAX HIGHT WHICH IS THE Y OF YOUR VERTEX

OpenStudy (anonymous):

t=0 is not a vertex

OpenStudy (anonymous):

that is 0 time e=meaning the apple hasnt been thrown yet...

OpenStudy (anonymous):

It is not a vertex because it is PART OF THE VERTEX. It is the x-intercept.

OpenStudy (anonymous):

how can your friend catch something that hasnt been thrown?

OpenStudy (anonymous):

@Nnesha Please confirm this.

OpenStudy (anonymous):

no x intercept is finding the root dude...

OpenStudy (anonymous):

you're really off

OpenStudy (anonymous):

@pooja195

OpenStudy (anonymous):

The question I posted before this was asking if a person threw an apple how many seconds would it take to reach the other person 3 stories up? We solved and got 2.4 seconds.

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

Show me how you will find the height then. You keep on complaining, but you're not showing any work.

OpenStudy (anonymous):

Of course it is correct. I was the one who helped her. :P

Nnesha (nnesha):

`(**HINT: The height would be the y-coordinate of the highest point.)`

OpenStudy (anonymous):

but now plug that number into the function

OpenStudy (anonymous):

that is the hight

OpenStudy (anonymous):

not t=0...

OpenStudy (anonymous):

@mathway im not calling you dumb people make mistakes and this is one of yours

OpenStudy (anonymous):

You would still get the same answer either way, 0.96.

OpenStudy (anonymous):

yes buy he cant say thats what it is you have to explain it in the context of the problem

OpenStudy (anonymous):

since its a root ofcourse it would cancle out the other 2 but in a 4th power polynomial that would fail miserably

OpenStudy (anonymous):

so the fundamental way to explain the answer is V = {(b/2a),f(b/2a)} and you look for Y

OpenStudy (anonymous):

Okay guys last question! Using the same equation - h(t) = -16t^2 + 38.4t + 0.96 With h = height and t = time in seconds 5) Suppose your friend does not catch the apple and it falls back to the ground. How long did it take the apple to hit the ground? (**HINT: When the apple hits the ground is when the graph hits the x-axis.)

OpenStudy (anonymous):

find the root

OpenStudy (anonymous):



Nnesha (nnesha):

calm plz use the hint highest point = max so we need y -coordinate of the max point (max=vertex)

Nnesha (nnesha):

lol sorry for late reply >.< internet iz very slow -.-

OpenStudy (anonymous):

lol its ok

OpenStudy (anonymous):

ok hold up go back what did you get for your sec for the apple to get to him??????!!!!

OpenStudy (anonymous):

the kid on the 3rd floor????

Nnesha (nnesha):

x coordinate of the vertex can be any number (0,y) = y-intercept :=)

OpenStudy (anonymous):

\[\frac{ -38.4 \pm \sqrt{38.4^{2}-4(-16)(0.96)} }{ 2(-16) }\]

OpenStudy (anonymous):

yea lol but only if it is centered on the orgin would that be a vertex

OpenStudy (anonymous):

yea @tbull1221 that works

OpenStudy (anonymous):

what does that give you?

OpenStudy (anonymous):

\[\frac{ -38.4 \pm \sqrt{1536} }{ -32 }\]

OpenStudy (anonymous):

Then you get: \[\frac{ -38.4 \pm 16\sqrt{6} }{ -32 }\]

OpenStudy (anonymous):

"x coordinate of the vertex can be any number (0,y) = y-intercept :=)" @Nnesha Thank you for backing up! At least I know that I'm not dumb and I don't necessarily have to say irrelevant things when I made a mistake. :)